The income statement for 2018 shows sale revenue of 400, depreciation of 50, net income of 50. If the dividend payout ratio was 20%, how much did retained earnings increase on the year-end balance sheet?
Answer #1

Retention ratio=1-dividend payout

=(1-0.2)=0.8

Hence increase in retained earnings=0.8*Net income

=(0.8*50)

which is equal to

=$40